OpenStudy (anonymous):

(4w^2-4w-8) - (2w^2+3w-6)

OpenStudy (texaschic101):

4w^2 - 4w - 8 - (2w^2 + 3w - 6) -- distribute the - through the parenthesis 4w^2 - 4w - 8 - 2w^2 - 3w + 6 now combine like terms....can you do that ?

OpenStudy (anonymous):

no show me!!

OpenStudy (texaschic101):

4w^2 - 2w^2 = ? -4w - 3w = ? -8 + 6 = ? can you answer these ?

OpenStudy (anonymous):

2w -7w -2??

OpenStudy (texaschic101):

close......2w^2 - 7w - 2

OpenStudy (texaschic101):

thats your answer

OpenStudy (texaschic101):

do you understand what like terms are ?

OpenStudy (anonymous):

yes thanks

OpenStudy (texaschic101):

your welcome :)